The booster club voted on where they would go for their annual trip. A majority of the club voted to go to a baseball game. They bought 29 tickets. Some of the tickets cost $21 each and some cost $27 each. The total cost of all the tickets was $675. How many tickets of each price did they buy?
Answer by Abbey(339)

let x = number of tickets @ $21
21x=amount of money these tickets cost
Let y= number of tickets @$27
27y = amount of money these tickets cost
the total number of tickets purchased:
x+y=29
The total value of the tickets purchased
21x+27y=$675
rewrite x+y=29 as
x=29-y
then substitute it into the value equation:
21(29-y)+27y=$675
clear the parentheses:
609-21y+27y=675
609+6y=675
subtract 609 from both sides:
6y=66
divide by 6
y=11
11 tickets were purchased at $27, so 18 were purchased at $29.
Check by multiplication
11*27=297
18*21=378
297+378=675